# A quasi-one-dimensional hydrogen-bonded monolayer ice phase

### v1

@@ III.2 Anomalous finite-temperature dependence of hydrogen bonding
Figure 2: a) The $\phi$ angle captures in-plane and out-of-plane fluctuations of water molecules’ dipole moments, while the $\theta$ angle captures in-plane rotations of water molecules. b) The left column contains the two-dimensional log probabilities along the $\phi$ and $\theta$ parameters for the flat-rhombic phase at 2 GPa and three different temperatures. The right column shows the average number of putative hydrogen bonds associated with each $(\theta,\phi)$ molecular orientation for the flat-rhombic phase under the same conditions. We employ the geometric definition of the hydrogen bond from Ref. 35.
@@
The characteristic interactions in the flat-rhombic phase lead to anomalous finite-temperature structure and hydrogen bonding dynamics. In bulk or conventionally hydrogen-bonded ice phases, such as the hexagonal and pentagonal monolayer phases, the most probable molecular orientations maximize the number of hydrogen bonds [25], and the protons maximize their residence time in orientations with the maximum possible hydrogen bonds. However, in the flat-rhombic phase, even at finite temperatures (extending even past 300 K), the most probable molecular orientations are not those that maximize hydrogen bonds but instead, those that maintain a distinct network of quasi-one-dimensional chains.
@@
To illustrate the flat-rhombic phases’ apparent “hesitancy” in forming additional hydrogen bonds, we computed the probabilities of molecular orientations defined by the angles $\theta$ and $\phi$, as presented in Fig. 2(a). These two-dimensional log probabilities are juxtaposed with the mean number of hydrogen bonds for each two-dimensional bin determined by the $\theta$ and $\phi$ angles, as depicted in Fig. 2(b). As shown in Appendix C, the nanoconfined hexagonal and pentagonal phases exhibit the expected trend: molecular orientations with more hydrogen bonds should be more stable. In contrast, the most probable flat-rhombic molecular orientations correspond to only two hydrogen bonds, even though orientations with three hydrogen bonds are possible. We begin to observe these orientations with three hydrogen bonds via the entropic exploration of $\theta$ and $\phi$ angles at high temperatures and in the presence of quantum nuclear fluctuations (see Appendix E). However, these orientations still do not emerge as local probability maxima, suggesting they are not stable or long-lived states. This behaviour reinforces the observation that the two-dimensional flat-rhombic ice phase prefers to remain hydrogen bonded only along one dimension, even at finite temperatures or in the presence of quantum nuclear fluctuations.
@@
Figure 3: a) The average number of hydrogen bonds across the range of temperatures in which the flat-rhombic phase is stable using classical and quantum simulations. b) The left plot shows the proportion of water molecules of each hydrogen bonding motif (defined in the text). The right plot shows that the 2D1A motif’s lifetime remains extremely short across the range of temperatures in which it is observed.
@@
III.3 Long-ranged spatial ordering and coherent proton dynamics
@@
Figure 4: a) The two symmetry-related structures of the flat-rhombic phase, with alternating rows of aligned dipoles, discerned by the different signs of $\sigma$. A zero value of $\sigma$ indicates that the molecular dipoles are not aligned row-wise. The side views in each panel show the system as viewed from the right side. b) For smaller unit cells with 144 water molecules, we observe coherent switching between these two states at intermediate temperatures. c) The free energy profiles for this $\sigma$ parameter show that raising the temperature lowers the free energy barrier between these states. The smooth, low-barrier free energy profile at high temperatures indicates that the molecular dipoles lose spatial coherence.
A consequence of the quasi-one-dimensional hydrogen bonded structure is that it may possess ferroelectric behaviour due to a net dipole moment along the hydrogen bonded chains. Ferroelectricity in ice has been conjectured previously in force field simulations [47], first-principles calculations of nanoconfined water [48], and experiments on confined and supported films [28, 49]. To investigate the dielectric behaviour of flat-rhombic ice, we model the temperature dependence of its dielectric response in Fig. 9 of Appendix D. Our results showed that the dielectric response of flat-rhombic ice remained small and consistent with the experiments of Ref. 50. The dielectric nature of flat-rhombic ice remains the same from 0 K up to its phase transition into a paraelectric hexatic (disordered) water phase [16]. Therefore, the potential ferroelectric dielectric response of flat-rhombic ice remains preserved even beyond room temperature, displaying the resilience of the quasi-one-dimensional structure to thermal and quantum nuclear fluctuations.
@@
Thus far, the results considered large simulation boxes with 576 molecules spanning tens of nanometers. However, in simulation boxes containing 144 molecules in a relatively small but experimentally accessible length scale of a few nanometers [27, 51], we identify unusual concerted dynamics at intermediate temperatures involving the breaking and forming of hydrogen bonds between equivalent structures. In this concerted motion, we observe an exchange between the individual molecular $\phi=\pm\pi/4$ states as a collective motion throughout the entire system, in which all the water molecules simultaneously swap the signs of their $\phi$ angles. To describe this motion as an exchange between these two states, we define a $\sigma$ order parameter in Appendix B. The structural change to the flat-rhombic phase between $\sigma$ values can be seen in Fig. 4(a).
@@
When the sign of $\sigma$ changes, the directions of the dipoles in each row switch in a concerted fashion, as seen in the side views of Fig. 4(a). During this motion, we also observe the complete rearrangement of the hydrogen bonding network in the lattice. Fig. 4(a) shows labelled water molecules in each $\sigma$ state, illustrating that the hydrogen bonds in these two states are between different pairs of water molecules. At low temperatures, the system remains frozen in a single $\sigma$ value. The resulting $\sigma$ free energy profile contains two minima separated by a large free energy barrier, shown in Fig. 4(c). At intermediate temperatures, we begin to see exchanges between $\sigma$ signs on the order of ${\sim}10$ ps, as shown in Fig. 4(b). The corresponding free energy profiles along $\sigma$ at various temperatures are shown in Fig. 4(c). Compared to the low-temperature setting, the intermediate temperature free energy profile exhibits a clear finite free energy barrier between these states. Eventually, at high enough temperatures, the free energy profile along $\sigma$ exhibits an extremely small free energy barrier, allowing the system to explore molecular configurations freely. We note that quantum nuclear motion makes the system more disordered due to zero-point fluctuations lowering the barrier for the system’s hydrogen bond network to switch between the free energy minima, as shown in Fig. 11 in Appendix E. Again, the observed length scale of concerted motion is of relevance to typical experimentally accessible or technologically relevant nanoscale systems extending to a few nanometers, such as small ice patches forming between graphene sheets [42] or high-pressure graphene nanobubbles formed by irradiation [51].
IV Conclusions
@@
We have analyzed the hydrogen bond structure and dynamics in the flat-rhombic phase of monolayer nanoconfined ice at first-principles-level accuracy using machine learning-driven simulations. Our work not only corroborates previous experimental and theoretical studies showing violations of ice rules at interfaces [52, 53, 26, 28] but also highlights an extreme scenario where hydrogen bonding plays a limited role in stabilizing the crystalline structure of ice. General ice rules must account carefully for the enthalpic penalty of breaking a hydrogen bond due to interactions with the confining walls and favourable vdW interactions. Instead of maximizing the number of hydrogen bonds, flat-rhombic ice forms one-dimensional chains stabilized by vdW interactions akin to quasi-one-dimensional vdW functional materials [32, 33]. This quasi-one-dimensional ice structure is robust with respect to thermal fluctuations and the rotational dynamics of water molecules, forming two or fewer hydrogen bonds up to its transition to the hexatic phase. We anticipate that direct evidence of the unique structure of the flat-rhombic ice can be experimentally investigated using sum-frequency generation spectroscopy [54] by disentangling the O–H vibrational stretching band [55] into donor-acceptor contributions.
@@
Like quasi-one-dimensional vdW functional materials that exhibit unique electronic, vibrational and optical properties [32, 33, 34], the flat-rhombic phase exhibits unusual properties with technological prospects. We observe long-ranged ordering of molecular dipoles along with a high spatial coherence in the hydrogen bond dynamics in the flat-rhombic ice phase, which is typically uncommon in systems extending to a few nanometers. This behaviour is consistent with the characteristics of a two-dimensional Ising model system. We anticipate that this concerted motion in flat-rhombic ice patches could be used to guide directional behaviours in nanoscale molecular devices that have already shown success in performing macroscopic-level tasks at surfaces by exploiting directed translational motion on short-length scales [56]. Coherent proton dynamics observed experimentally in bulk systems, such as proton tunnelling in hexagonal ice [57] and dielectric phase transitions in molecular ferroelectrics [58, 59], have already been computationally detected on much smaller scales [60, 61, 62] in flat-rhombic ice. This suggests that confinement effects might be crucial in phenomena like tunnelling-induced or dielectric phase transitions. Our work opens doors to exploring new confined molecular materials with the structure-function relationships of low-dimensional functional materials that exhibit technologically relevant properties.

### v2

@@ II.2 Anomalous finite-temperature dependence of hydrogen bonding
Fig. 3: |||| Molecular orientation and the number of putative hydrogen bonds. a) The two-dimensional log probabilities along the $\phi$ and $\theta$ parameters for the flat-rhombic phase at 2 GPa and three different temperatures. b) The average number of putative hydrogen bonds associated with each $(\theta,\phi)$ molecular orientation for the flat-rhombic phase under the same conditions. Relevant molecular orientations are depicted in Fig. 2(c). We employ the geometric definition of the hydrogen bond from Ref. 29. Source data are provided as a Source Data file.
@@
The characteristic interactions in the flat-rhombic phase lead to anomalous finite-temperature hydrogen bonding structure and dynamics. In bulk or conventionally hydrogen-bonded ice phases, such as the hexagonal and pentagonal monolayer phases, the most probable molecular orientations are the ones that maximize the number of hydrogen bonds [25], and the protons maximize their residence time in orientations with the maximum possible hydrogen bonds. However, in the flat-rhombic phase, even at finite temperatures (extending past 300 K), the most probable molecular orientations are not those that maximize hydrogen bonds but instead, those that maintain a distinct network of quasi-one-dimensional chains.
To characterize the rotational motion of individual water molecules, we define $\phi$ and $\theta$ angles for molecular dipoles. These are depicted in Fig. 2(a) and  2(b), respectively. These figures also indicate the coordinate axes used to define these angles. The $\phi=\frac{\pi}{2}-\text{arccos}\left(\mathbf{\hat{\bm{\mu}}}\cdot\mathbf{\hat{z}}\right)$ angle captures the alignment between the molecular dipole vector $\bm{\mu}$ and the $z$-axis, with hats above vectors indicating normalized unit vectors. A water molecule with $\phi=\pm\;\pi/2$ will have its dipole vector aligned with the $\pm$ $z$-axis. We also wrap around the $\phi$ angles computed such that $\phi$ always lies in the range $[-\frac{\pi}{2},+\frac{\pi}{2}]$. The $\theta=\text{arctan}\left({\mathbf{c}\cdot\mathbf{\hat{z}}}/{\mathbf{c}\cdot\mathbf{\hat{x}}}\right)$ angle captures a molecule’s rotation about its center of mass around the $y$ axis with $\mathbf{c}=\mathbf{r_{\text{OH}_{\text{1}}}}\times\mathbf{r_{\text{OH}_{\text{2}}}}$ and $\mathbf{r_{\text{OH}_{\text{1}}}}$ and $\mathbf{r_{\text{OH}_{\text{2}}}}$ being the OH bond vectors of an individual water molecule. $\theta$ measures the rotation of $\mathbf{c}$ around the $y$-axis. Since the choice of $\text{H}_{1}$ and $\text{H}_{2}$ is arbitrary, we wrap around the $\theta$ angles so that they always lie in the range $[0,\pi]$.
@@
To illustrate the flat-rhombic phases’ apparent hesitancy to form additional hydrogen bonds, we computed the probabilities of molecular orientations defined by these $\theta$ and $\phi$ angles. These two-dimensional log probabilities are juxtaposed with the mean number of hydrogen bonds for each two-dimensional bin determined by the $\theta$ and $\phi$ angles, as depicted in Fig. 3. As shown in Supplementary Note V, the nanoconfined hexagonal and pentagonal phases exhibit the expected trend: molecular orientations with more hydrogen bonds should be more stable. In contrast, the most probable flat-rhombic molecular orientations (depicted in the left and right plots of Fig. 2(c)) correspond to only two hydrogen bonds. This is true even though orientations with three hydrogen bonds (like the one in the middle plot of Fig. 2(c)) are possible. We begin to observe these orientations with three hydrogen bonds via the entropic exploration of $\theta$ and $\phi$ angles at high temperatures and in the presence of quantum nuclear fluctuations (see Supplementary Note VI). However, these orientations still do not emerge as local probability maxima, suggesting they are not stable or long-lived states. This behavior reinforces the observation that the two-dimensional flat-rhombic ice phase prefers to remain hydrogen bonded only along one dimension, even at finite temperatures.
@@
Fig. 4: |||| Temperature dependence of hydrogen bonding. a) The average number of putative/geometric hydrogen bonds across the range of temperatures in which the flat-rhombic phase is stable using classical and quantum simulations. The error bars show the standard error of the mean, as computed by block averaging over 10 blocks. b) The proportion of water molecules of each hydrogen bonding motif (defined in the text). The 1D2A proportions are perfectly aligned with the 2D1A proportions. c) The 2D1A motif’s lifetime remains extremely short across the range of temperatures in which it is observed. The error bars show the standard error of the mean, computed across all instances of hydrogen bonding. Dashed lines serve as a visual guide for the eye. Source data are provided as a Source Data file.
@@
This apparent increase emerges from the shortcoming of the standard geometric definition of a hydrogen bond, which cannot distinguish fluctuating molecular orientations from long-lived orientations that correspond to true hydrogen bonds. As shown in Fig. 4(c), the lifetime of the 2D1A orientations is extremely short – on the order of femtoseconds. As suggested by Schienbein and Marx [30], these fleeting hydrogen bonds are best referred to as either putative or geometric, as they do not last long enough to be considered hydrogen bonds from a traditional perspective. In Supplementary Note II, we compute the number of hydrogen bonds that survive for typical intermolecular oscillations. Within this definition by Schienbein and Marx [30], a hydrogen bond is not counted if its lifetime is shorter than the period of typical intermolecular oscillations. The resulting dynamical hydrogen bond count indeed exhibits the expected disorder-induced decrease as we raise temperature.
II.3 Dielectric nature of the flat-rhombic phase
Fig. 5: |||| Temperature dependence of the dielectric response. The classical dielectric constant of the flat-rhombic phase across the range of temperatures considered in this work. The singularity at 380 K corresponds to the transition to the hexatic phase of nanoconfined water [15]. Source data are provided as a Source Data file.
A consequence of the quasi-one-dimensional hydrogen-bonded structure is that it may possess ferroelectric behavior due to its net dipole moment along the direction of the hydrogen-bonded chains. Ferroelectricity in ice has been conjectured previously in force field simulations [2], first-principles calculations of nanoconfined water [43], and experiments on confined [28] and supported [44] films. To investigate the dielectric behavior of flat-rhombic ice, we model the temperature dependence of its dielectric response in Fig. 5. We analyzed the in-plane $\varepsilon_{\parallel}$ and out-of-plane $\varepsilon_{\perp}$ dielectric constants based on the variance of the system’s polarization. Since our MLP only predicts the potential energy surfaces, a first principles investigation of the dielectric response would be computationally demanding due to numerous single point calculations of the electronic polarization [45]. Therefore, for a semi-quantitative understanding, we use a simple linear polarization model based on TIP4P charges [33].
We select the TIP4P water model due to its low computational cost and a semi-quantitative description of the dielectric response of bulk [33] and confined [47] water. The TIP4P model underpredicts the polarization of aqueous systems as it doesn’t incorporate the electronic polarization of the water molecules. For instance, TIP4P predicts a molecular dipole moment of 2.348 D which is 12.8% lower compared to first principles estimates [48], and effectively leads to a ${\sim}25$% lower dielectric response. To incorporate the lack of the electronic polarization in the TIP4P model, we rescale the calculated dielectric response by a factor of 1.25, as has been done previously [47]. We report the classical dielectric response, as quantum nuclear effects are expected to only make a small quantitative difference [33].
We observe low and near-constant in-plane and out-of-plane dielectric constants for the flat-rhombic phase, in agreement with the experiments on nanoconfined water from Ref. 49. The low dielectric response of this phase is only an indirect outcome of vdW interactions, as they dictate the thermodynamic stability of the flat rhombic phase. The dielectric nature of flat-rhombic ice remains the same from 0 K up to its phase transition into the paraelectric hexatic (disordered) water phase [8]. The temperature independence of the flat-rhombic phase’s dielectric behaviors reflects the resilience of the quasi-one-dimensional structure to thermal and quantum nuclear fluctuations. The singularity observed in Fig. 5 indicates the temperature at which the phase transition to the hexatic phase occurs. The hexatic phase exhibits an increased in-plane dielectric constant but a similarly small out-of-plane dielectric constant.
II.4 Long-ranged spatial ordering and coherent proton dynamics
@@
Fig. 6: |||| Hydrogen bonded switching behavior. a) The two symmetry-related structures of the flat-rhombic phase, with alternating rows of aligned dipoles, discerned by the different signs of $\sigma$. A zero value of $\sigma$ indicates that the molecular dipoles are not aligned row-wise. The side views in each panel show the system as viewed from the right side, where the switch in the row-dependent dipole direction between the two states is clear. b) For smaller unit cells with 144 water molecules, we observe coherent switching between these two states at intermediate temperatures, as shown in the $\sigma$ trajectories for three different temperatures. c) The free energy profiles for this $\sigma$ parameter show that raising the temperature lowers the free energy barrier between these states. The smooth, low-barrier free energy profile at high temperatures indicates that the molecular dipoles lose spatial coherence. Source data are provided as a Source Data file.
@@
Thus far, the results considered large simulation boxes with 576 molecules spanning tens of nanometers. However, in simulation boxes containing 144 molecules in a relatively small but experimentally accessible system size on the order of nanometers [18, 51], we identify unusual concerted dynamics at intermediate temperatures involving the breaking and forming of hydrogen bonds between equivalent structures. In this concerted motion, we observe an exchange between the individual molecular $\phi=\pm\pi/4$ states as a collective motion throughout the entire system, in which all the water molecules simultaneously swap the signs of their $\phi$ angles (see Supplementary Video 1).
@@
To describe this concerted motion, we define a $\sigma$ order parameter (see supporting information section for details) to distinguish the two states shown in Fig. 6(a). The structural change to the flat-rhombic phase between $\sigma$ values can be seen in Fig. 6(a). When the sign of $\sigma$ changes, the directions of the dipoles in each row switch in a concerted fashion. During this motion, we also observe the complete rearrangement of the hydrogen bonding network in the lattice. Fig. 6(a) shows labeled water molecules in each $\sigma$ state, illustrating that the hydrogen bonds in these two states are between different pairs of water molecules. At low temperatures, the system remains frozen in a single $\sigma$ value. The resulting $\sigma$ free energy profile contains two minima separated by a large free energy barrier, shown in Fig. 6(c). At intermediate temperatures, we begin to see exchanges between $\sigma$ signs on the order of ${\sim}10$ ps, as shown in Fig. 6(b). Compared to the low-temperature setting, the intermediate temperature free energy profile exhibits a clear finite free energy barrier between these states. Eventually, at high enough temperatures, the free energy profile along $\sigma$ exhibits an extremely small free energy barrier, allowing the system to explore molecular configurations freely. While the highest temperatures explored here are greater than the melting temperature of ice, we believe that our trajectories correspond to a metastable state associated with the solid phase. Since the free energy profiles at the intermediate and high-temperature conditions are qualitatively the same, the near-free exploration of molecular configurations in the high-temperature simulations is a result of thermal activation. To confirm this hypothesis, we compute rotational autocorrelation functions at these different temperatures in Supplementary Note IV.
We also note that quantum nuclear motion makes the system more disordered due to zero-point fluctuations lowering the barrier for the system’s hydrogen bond network to switch between the free energy minima, as shown in Supplementary Figure 8. Furthermore, since nuclear quantum effects lead to increased proton disorder, the $\sigma$ parameter gets pushed closer to 0 on average. This causes the minima in the free energy profiles along $\sigma$ to move closer to 0 in our PIMD simulations as compared to the free energy profiles from our classical simulations.
As stated earlier, this concerted motion is not observed in the larger simulation cell size that contains 576 molecules for the timescales we consider: it is only seen in simulations with 144 water molecules. This suggests that the free energy barrier for this concerted motion is size-extensive. As a result, we would expect a larger system size to either exhibit an ordered phase or a phase in which smaller domains that exhibit this behavior coexist. However, crystallites of nanoconfined water molecules containing around 100 water molecules are experimentally accessible, so such concerted motion could be observed in experiments as small ice crystallites forming between graphene sheets [18] or high-pressure graphene nanobubbles formed by irradiation [51]. Previous work has suggested that the simulation setup we employ is still reasonable for studying such encapsulated systems, even without explicit confining atoms [18, 52]. A careful treatment of such encapsulated systems would require a detailed analysis of the water-carbon interactions at the edge of the confined water pockets. In this work, we ignore these edge effects and instead focus on characterizing the behavior of nanoconfined water under idealized atomistically flat nanoconfinement. A careful consideration of the edge effect would be an interesting and relevant topic for future work.
III Discussion
@@
We have analyzed the hydrogen bond structure and dynamics in the flat-rhombic phase of monolayer nanoconfined ice at first-principles-level accuracy using machine learning-driven simulations. Our work not only corroborates previous experimental and theoretical studies showing violations of ice rules at interfaces [7, 14, 53, 26, 28] but also highlights an extreme scenario where hydrogen bonding plays a limited role in stabilizing the crystalline structure of ice. General ice rules must account carefully for the enthalpic penalty of breaking a hydrogen bond due to interactions with the confining walls and favourable vdW interactions. Instead of maximizing the number of hydrogen bonds, flat-rhombic ice forms one-dimensional chains stabilized by vdW interactions akin to quasi-one-dimensional vdW functional materials [32, 33]. The ZZ-qBI phase identified in Ref. 14 is stable at higher lateral pressures than the flat-rhombic phase, yet it exhibits a similar behavior, indicating that such hydrogen bond topologies may occur across a broad range of nanoconfined conditions. This quasi-one-dimensional ice structure is robust with respect to thermal fluctuations and the rotational dynamics of water molecules, forming two or fewer hydrogen bonds up to its transition to the hexatic phase. We anticipate that direct evidence of the unique structure of the flat-rhombic ice can be experimentally investigated using sum-frequency generation spectroscopy [54] by disentangling the O–H vibrational stretching band [15] into donor-acceptor contributions.
@@
Like quasi-one-dimensional vdW functional materials that exhibit unique electronic, vibrational, and optical properties [32, 33, 34], the flat-rhombic phase exhibits unusual properties with technological prospects. We observe long-ranged ordering of molecular dipoles along with a high spatial coherence in the hydrogen bond dynamics in the flat-rhombic ice phase, which is typically uncommon in systems containing on the order of 100s of molecules. This behavior is consistent with the characteristics of a two-dimensional Ising model system. We anticipate that this concerted motion in flat-rhombic ice patches could be used to guide directional behaviors in nanoscale molecular devices that have already shown success in performing macroscopic-level tasks at surfaces by exploiting directed translational motion on the nanometer scale [55]. Coherent proton dynamics observed experimentally in bulk systems, such as proton tunneling in hexagonal ice [56] and dielectric phase transitions in molecular ferroelectrics [57, 58], have already been computationally detected on much smaller scales [59, 60, 61] in flat-rhombic ice. This suggests that confinement effects might be crucial in phenomena like tunneling-induced or dielectric phase transitions. Our work opens doors to exploring new confined molecular materials with the structure-function relationships of low-dimensional functional materials that exhibit technologically relevant properties.
IV Methods
Confining Potential
The nanoconfined system considered in this work is a single layer of water molecules trapped between two parallel sheets – mimicking the experimental setup in Ref. 27. We model the interactions between the sheets and the water molecules using a simple Morse potential fit to water-carbon QMC interaction energies [27]. Such a potential, characterized by perfectly smooth walls, has been widely adopted in previous studies, both in force field [9, 6, 5] and first-principles research [3, 4, 9, 7, 8, 14]. The uniform confinement potential has demonstrated semi-quantitative accuracy in describing the behavior of water confined within graphene-like cavities [6], as corroborated by a good agreement of stable phases and melting temperatures with respect to confinement simulations that include explicit carbon atoms [52]. Furthermore, the uniform confinement model’s atomistically flat nature allows for a clean interpretation of topological confinement effects – a phenomenon that extends beyond the specific context of graphene-based confinement.
Machine Learning Potential
For water-water interactions, we employ a newly-trained MLP [10, 13] at the revPBE0-D3 level [11, 12] – an appropriate dispersion-corrected hybrid-functional DFT level for bulk, interfacial and nanoconfined water [66, 67, 15, 8]. We report the training and validation protocols for the MLP in Supplementary Note IB. Our simulations are run in the $NP_{xy}T$ ensemble, where $P_{xy}$ denotes the lateral pressure, or vdW pressure [18], that acts in the $x$ and $y$ directions. This lateral pressure emulates the net lateral forces on the water molecules due to the edges of the confining pocket (further details on the origins of this lateral pressure can be found in Supplementary Note IC). Both simulations [52] and experiments [18] have estimated this vdW pressure to be on the gigapascal scale. Hence, our simulations are run at a lateral pressure of 2.0 GPa across various temperatures within the metastability range of the flat-rhombic phase [8]. We refer the reader to Supplementary Note I for more details on the simulations
First-Principles Calculations
To estimate the (zero temperature) static stabilization energy of a nanoconfined ice crystal or a quasi-one-dimensional chain of water molecules, we perform direct first-principles calculations using the CP2K code [68] employing the revPBE0-D3 functional and the convergence parameters from Ref. 8. We perform these calculations on geometry-optimized structures of the confined ice phases [14]. The geometry-optimized structures of the hexagonal, pentagonal, and flat-rhombic monolayer phases were taken from Ref. 8. In contrast, the structure of the zigzag quasi bilayer ice from Ref. 14 was optimized using the CP2K code [68]. We define the stabilization energy of a crystal with respect to an isolated molecule as $E_{\textrm{lattice}}=E_{\textrm{crystal}}-E_{\textrm{gas}}$, where $E_{\textrm{crystal}}$ is the single-point energy of the crystalline lattice per water molecule, and $E_{\textrm{gas}}$ is the single-point energy of an isolated water molecule in vacuum. Similarly, we define the stabilization energy of a chain of water molecules with respect to an isolated molecule as $E_{\textrm{chain}}=E_{\textrm{chain}}^{\textrm{struct.}}-E_{\textrm{gas}}$, where $E_{\textrm{chain}}^{\textrm{struct.}}$ is the single-point energy of the chain structure, and the stabilization energy of a crystal with respect to a chain of water molecules as $E_{\textrm{stack}}=E_{\textrm{lattice}}-E_{\textrm{chain}}$. We also estimate these quantities without the D3 dispersion correction to assess the role of vdW interactions.
Data Availability
The data supporting the findings of this study are openly available on GitHub (https://github.com/water-ice-group/hbond_nanoconfined_water). This includes all simulation scripts and starting structures. Source data are provided with this paper. They are also available via this GitHub repository.
Code Availability
All of the analysis scripts supporting the findings of this study are openly available on GitHub (https://github.com/water-ice-group/hbond_nanoconfined_water).